# TPFL: Tsetlin-Personalized Federated Learning with Confidence-Based Clustering

### v2

@@ 5.4 Phase D
1:Input: Received weight vectors $W_{m}^{{c}_{\text{max}}}$ and its corresponding class $c_{max}$ from client $m$, local dataset $D_{\text{test}}$.
2:Output: Aggregated weight vector $W_{\text{cluster}}^{k}$ for each cluster $k$.
3:for each client $m$ do
4:     $k\leftarrow c_{max}$triangleright class $c_{max}$ determines cluster $k$ for $W_{m}$.
5:     if weights $W_{m}^{{c}_{\text{max}}}$ are initial weights in cluster $k$ then
6:         Initialize cluster weights $W_{\text{cluster}}^{k}$ for each cluster $k$:
@@
7:     else
8:         Aggregate received weights:
@@
9:     end if
10:end for
11:Compute average of weights in the cluster:
@@
12:Send updated weight vectors $W_{\text{cluster}}^{k}$ to corresponding clients of each cluster
13:Evaluate clients on $D_{\text{test}}$
@@ 6.6 Baselines
    FedAvg: This was the pioneering method that is still considered the most basic and one of the most important baselines in the FL research community [32].
@@ 6.7 TPFL Results
This evident pattern clearly proves the relationship between the data distribution of all clients and their confidence. In other words, the more skewed the data distribution, the more confident the model towards that class, and the the more confident the model, the faster their convergence time. Hence, the convergence time for our confidence-based clustering method is far less when the clients’ partitioning scheme is severely non-IID. On the other hand, the models tend to be less confident when data distribution is uniformly IID, which is a clear indication that the entire federation requires much more time in order to become more confident and reach high-accuracy results. For such circumstances, increasing the local clients’ training epochs or even increasing the total round number of federation can be beneficial so that high-accuracy results can be achieved. Figure 5 shows the communication cost of all three datasets under 5 different experimental setups for a total of 10 rounds. The download cost is far lower than the upload cost, however, the more classes the dataset has, the more probable it is for the download cost to be more. Therefore, it can be seen for the FEMNIST dataset under experiment 5 that the download cost progressively becomes closer to the upload cost
@@ 6.8 Comparison
Clients in the FL settings are normally evaluated under non-IID data partitioning schemes. Therefore, We compared the baseline methods with the proposed TPFL method under non-IID data partitioning. Hence, our comparison is between the best results of each baseline with non-IID data partitioning and TPFL under experiment 5. Table 6.7 shows the summary of our comparison. As it can be seen, the TPFL model outperforms all baselines in all datasets followed by IFCA and then FedTM. It is worth mentioning that TFPL implementation was carried out only in 10 rounds, the least number of rounds among all baseline methods. This phenomenon illustrates the positive effect of cluster-based personalization within TPFL setting. In addition, the communication cost of the TPFL in terms of uploaded and downloaded model parameters is the most optimized followed by FedTM, which itself is based on the TM algorithm. This evident result shows the difference between DL-based models and the TM-based models. Based on the obtained results and their comparison with baseline models, it can be understood that the TPFL approach can be utilized in a one-shot learning fashion, in which each experimental setup is carried out only in one round. However, it is advisable to increase the number of local training epochs in order to make up for absence of multiple rounds of training.
@@ 7 Conclusion
With the advent of FL algorithms, numerous works have been put forward to maintain the users privacy However, vanilla FL suffers from numerous disadvantages. PFL is considered a promising solution to solve the shortcoming of the vanilla FL. In this work, we propose a personalization method for the TM algorithm based on confidence-based clustering, which we dub Tsetlin-Personalized Federated Learning (TPFL). The results of TPFL compared with other FL and PFL baselines demonstrate that TPFL outperforms all baselines for MNIST, FashionMNIST and FEMNIST dataset. A promising future work that can be built upon TPFL is the utilization of innovative clustering strategies such as thresholding strategies for defining a confidence threshold and multiclass selection of class weights - sharing weights of more than one class.

### v3

@@ 5.4 Phase D
Input: Received weight vectors $W_{m}^{{c}_{\text{max}}}$ and its corresponding class $c_{max}$ from client $m$, local dataset $D_{\text{test}}$.
Output: Aggregated weight vector $W_{\text{cluster}}^{k}$ for each cluster $k$.
1:for each client $m$ do
2:     $k\leftarrow c_{max}$triangleright class $c_{max}$ determines cluster $k$ for $W_{m}$
3:     if weights $W_{m}^{{c}_{\text{max}}}$ are initial weights in cluster $k$ then
4:         Initialize cluster weights $W_{\text{cluster}}^{k}$ for cluster $k$:
@@
5:     else
6:         Aggregate received weights:
@@
7:     end if
8:end for
9:Compute average of weights in the cluster:
@@
10:Send updated weight vectors $W_{\text{cluster}}^{k}$ to corresponding clients of each cluster
11:Evaluate clients on $D_{\text{test}}$
@@ 6.6 Baselines
    FedAvg: This was the pioneering method that is still considered one of the most important baselines in the FL research community [33].
@@ 6.7 TPFL Results
This evident pattern clearly proves the relationship between the data distribution of all clients and their confidence. In other words, the more skewed the data distribution, the more confident the model towards that class, and the more confident the model, the faster their convergence time. Hence, the convergence time for our confidence-based clustering method is far less when the clients’ partitioning scheme is severely non-IID. On the other hand, the models tend to be less confident when data distribution is uniformly IID, which is a clear indication that the entire federation requires much more time in order to become more confident and reach high-accuracy results. For such circumstances, increasing the local clients’ training epochs or even increasing the total round number of federation can be beneficial so that high-accuracy results can be achieved. Figure 5 shows the communication cost of all three datasets under 5 different experimental setups for a total of 10 rounds. The download cost is far lower than the upload cost, however, the more classes the dataset has, the more probable it is for the download cost to be more. Therefore, it can be seen for the FEMNIST dataset under experiment 5 that the download cost progressively becomes closer to the upload cost
@@ 6.8 Comparison
Clients in the FL settings are normally evaluated under non-IID data partitioning schemes. Therefore, We compared the baseline methods with the proposed TPFL method under non-IID data partitioning. Hence, our comparison is between the best results of each baseline with non-IID data partitioning and TPFL under experiment 5. Table 6.7 shows the summary of our comparison. As it can be seen, the TPFL model outperforms all baselines in all datasets followed by IFCA and then FedTM. It is worth mentioning that TFPL implementation was carried out only in 10 rounds, the least number of rounds among all baseline methods. This phenomenon illustrates the positive effect of cluster-based personalization within TPFL setting. In addition, the communication cost of the TPFL in terms of uploaded and downloaded model parameters is the most optimized followed by FedTM, which itself is based on the TM algorithm.
This evident result shows the difference between DL-based models and the TM-based models. Based on the obtained results and their comparison with baseline models, it can be understood that the TPFL approach can be utilized in a one-shot learning fashion, in which each experimental setup is carried out only in one round. However, it is advisable to increase the number of local training epochs in order to make up for absence of multiple rounds of training.
@@ 7 Conclusion
With the advent of FL algorithms, numerous works have been put forward to maintain the users privacy. However, vanilla FL suffers from numerous disadvantages. PFL is considered a promising solution to solve the shortcomings of the vanilla FL. In this work, we propose a personalization method called Tsetlin-Personalized Federated Learning (TPFL) that leverages the voting mechanism of the TM in order to cluster clients based on their confidence towards a class. The results of TPFL compared with other FL and PFL baselines demonstrate that TPFL outperforms all baselines for MNIST, FashionMNIST and FEMNIST datasets. A promising future work that can be built upon TPFL is the utilization of innovative clustering strategies such as thresholding strategies for defining a confidence threshold and multiclass selection of class weights | sharing weights of more than one class | so that a client can be assigned to multiple clusters.